OpenStudy (anonymous):

will someone explain to me why -2 1/3 +3 1/4= 11/12

OpenStudy (jdoe0001):

hmm what's there to explain?

OpenStudy (anonymous):

when you make -2 1/3 why is it -7/3 and not -6?

OpenStudy (anonymous):

*-6/3

OpenStudy (anonymous):

lol that question didn't make sense

OpenStudy (jdoe0001):

the negative IN FRONT of the fraction applies NOT just to the 2 after it, but to the whole fraction \(-2\frac{1}{3}+3\frac{1}{4}=\cfrac{11}{12}\qquad \begin{cases} -2\frac{1}{3}\implies \cfrac{-(2\cdot 3+1)}{3}\implies \cfrac{-7}{3}\\ 3\frac{1}{4}\implies \cfrac{3\cdot 4+1}{4}\implies \cfrac{13}{4} \end{cases}\)

OpenStudy (anonymous):

oh ok thank you!

OpenStudy (jdoe0001):

\(-2\frac{1}{3}\iff -\left( 2\frac{1}{3} \right)\)

OpenStudy (jdoe0001):

yw